OpenStudy (anonymous):

evaluate x^2-3y+x when x=9 and y=7

OpenStudy (bibby):

Plug in the given values in to the variables and evaluate. Can you solve?\[9^2-3(7)+9=?\]

OpenStudy (anonymous):

Yup is it 81-21+9

OpenStudy (bibby):

exactly. And you can evaluate it however you feel comfortable. For example 81+9-21 = 90 -21 (81-21)+9=60+9

OpenStudy (anonymous):

Oh ok thanks

OpenStudy (anonymous):

-4(x+6)+5=6(x+2)+5